Compute the electric potential from a source charge of -2.33 x

10^-13 C if a test charge will be placed 6.45 x 10^−15 m from it.


1
Expert's answer
2021-10-20T10:42:28-0400

By the definition of the electric potential, we have:


"V=\dfrac{kQ}{r}=\dfrac{9.0\times10^{9}\ \dfrac{N\times m^2}{C^2}\times(-2.33\times10^{-13}\ C)}{6.45\times10^{-15}\ m}=-3.25\times10^{11}\ V."